# S3-Attention:Attention-Aligned Endogenous Retrieval for Memory-Bounded Long-Context Inference

### v1

@@ Abstract
Large language models are increasingly used on multi-document and long-form inputs, yet long-context inference remains memory- and noise-inefficient: KV caching scales linearly with context length, and external retrievers may return lexically similar but causally irrelevant passages. We present S3-Attention, a memory-first inference-time framework that treats long-context usage as attention-aligned endogenous retrieval. S3-Attention decodes transient key/query projections into Top-k sparse feature IDs via lightweight sparse autoencoders and builds a CPU inverted index (feature → token positions/spans) during a streaming scan, allowing the KV cache to be discarded and keeping GPU memory bounded by the scan chunk size. At generation time, feature co-activation retrieves compact evidence spans, optionally fused with BM25 for exact matching. Under a unified LongBench evaluation protocol (fixed prompting/decoding and matched token budgets), S3-Hybrid closely matches full-context inference on multiple model families and improves robustness on several information-dense settings. We also report the current engineering limitation that our prototype incurs higher wall-clock latency than optimized FullKV baselines, motivating future kernel-level optimization.
@@ 1 Introduction
Large language models (LLMs) have become a default interface for question answering, summarization, and multi-step reasoning. As these systems move from single documents to collections of reports, codebases, and conversational histories, they increasingly need to condition on contexts that far exceed a single attention window(Ding et al., 2024; Zhao et al., 2024a). The prevailing response is to scale context lengths through continued training or architectural modifications(Liu et al., 2025b; Chen et al., 2024b; Hu et al., 2024; Du et al., 2025; Mao et al., 2025; Mohtashami and Jaggi, 2023), with recent models advertising windows from 128K up to million-token regimes.
@@
However, longer context windows do not automatically translate into reliable long-context reasoning. Full-context inference remains deployment-unfriendly(Ma et al., 2024): self-attention incurs quadratic compute(Lou et al., 2024), and KV caches scale(Sun et al., 2024) linearly with the number of processed tokens(Tang et al., 2024), quickly saturating GPU memory(Wang et al., 2025a). At the same time, real-world long inputs are typically sparse in signal—only(Zhu et al., 2025) a small fraction of tokens are causally useful for answering a given query—so naively attending to everything can dilute evidence and amplify distraction(Xu et al., 2024; Hooper et al., 2025; Zarch et al., 2025).
A pragmatic alternative is Retrieval-Augmented Generation (RAG)(Lewis et al., 2020; Cheng et al., 2025), which replaces full-context attention with an external retriever that selects a small set of passages for the generator. While RAG is memory-efficient and widely used, it introduces a semantic mismatch: retrieval is performed in an embedding space whose similarity structure need not align with the generator’s internal reasoning features(Wei et al., 2025). In noisy or heterogeneous corpora, this mismatch can surface lexically overlapping yet causally irrelevant text, degrading multi-hop reasoning(Liu et al., 2025a) and increasing hallucinations(Wang et al., 2025b).
These limitations motivate a complementary direction: endogenous retrieval, where the model retrieves evidence using its own internal signals(Wu et al., 2024). Recent results support this premise(Zhang et al., 2025). For instance, InfiniRetri links attention allocation patterns to retrieval behavior and leverages attention information to retrieve needles from extremely long inputs without additional training (Ye et al., 2025). Meanwhile, KV-cache compression and sparse-attention methods use attention scores(Xiao et al., 2024a) or other internal statistics to decide which past states to retain. Yet directly operating on dense attention weights or cached states does not provide a scalable, queryable memory(Chen et al., 2025b): the representations are continuous, high-dimensional, and expensive to store or search at token granularity(Ma et al., 2024; Liu et al., 2024, 2025c), and most systems are optimized for efficiency rather than explicitly surfacing interpretable evidence spans.
@@
We propose S3-Attention (Sparse & Semantic Streaming Attention), which turns long-context inference into a streaming, attention-aligned retrieval procedure. The key idea is to compress the transient key and query projections—already computed by the model—into discrete sparse semantic features that can be indexed efficiently. Concretely, we train lightweight Top-$k$ sparse autoencoders (SAEs) on attention projections and use them at inference time to map each token span to a small set of active features. During a single streaming prefill pass, S3-Attention encodes and stores an inverted index (feature $\rightarrow$ spans) on CPU and then discards the KV cache, achieving O(1) GPU memory with respect to total context length. At generation time, the query’s SAE features retrieve high-density spans via feature co-activation, which we refine with density smoothing and non-maximum suppression. To improve robustness on exact entity matching and other lexical cues, we optionally fuse the endogenous score with BM25, yielding S3-Hybrid.
@@
We evaluate S3-Attention on LongBench across multiple LLM families and tasks in a unified inference environment. S3-Hybrid retains 99.4% of full-context performance on Llama-3-8B (24.87 vs. 25.01) and over 99% on Qwen2-7B (18.80 vs. 18.92), while enabling constant-GPU-memory long-context processing. Beyond average scores, we observe consistent gains on multi-document and multi-hop settings and a “denoising” effect on information-dense tasks where selectively filtering context can outperform full-context attention.
@@
  • We articulate long-context inference as an endogenous retrieval problem and motivate why aligning retrieval with internal attention representations can mitigate the semantic gap of exogenous retrievers.
  • We introduce S3-Attention, which uses SAE-decoded sparse semantic features to build a streaming inverted index and achieve O(1) GPU memory without fine-tuning the base LLM.
  • We demonstrate strong empirical performance on LongBench and show that fusing endogenous semantic signals with BM25 yields a robust hybrid retriever with near-lossless fidelity to full-context inference.
S3-Attention is designed for settings where GPU memory is the primary bottleneck and only a small fraction of a long input is causally useful for a given query. Our goal is not to outperform all KV-compression or advanced RAG pipelines in wall-clock latency, but to provide a queryable, attention-aligned indexing mechanism that (i) avoids storing dense KV states, (ii) retrieves interpretable evidence spans, and (iii) can be combined with lexical retrieval for robustness.
@@ 3.2 Theoretical Framework: The Endogenous Information Bottleneck
Instead, we propose an operational endogenous proxy derived from the model’s own attention-matching signals. Specifically, we discretize transient attention projections into a sparse feature space and rank context positions by their inverse-document-frequency (IDF) weighted feature co-activation with the query. Appendix B provides a motivating inequality under explicit simplifying assumptions. We emphasize that this result should be interpreted as a heuristic justification of our scoring rule, rather than as a guarantee that feature overlap faithfully tracks mutual information in real models.
@@ 4.2 Qualitative Analysis: Mechanism of Endogenous Alignment
Figure 2: Endogenous vs. Exogenous Retrieval. Top: RAG (BGE-Small) is distracted by high lexical overlap, ranking a generic biography (Sentence 1) higher than the true answer (Sentence 5). Bottom: S3-Attention (Ours) ignores the distraction, showing sparse activation peaks exclusively at the semantic answer anchor (“The Post”) and its reasoning evidence (“Pentagon Papers”).
Additional visualization results and analysis of other datasets are provided in Appendix F.
4.3 Quantitative Results on LongBench
We present the comprehensive evaluation results on LongBench in Table 3. It is important to note that absolute evaluation scores on LongBench can fluctuate significantly depending on inference environments (e.g., prompt templates and quantization kernels). To ensure a rigorous comparison, we distinguish between standard baselines (provided for reference) and methods evaluated within our unified environment (marked in bold), which include FullKV, RAG, BM25, and our S3-Attention variants.
Performance Retention: Near-Lossless Compression.
  • Superior Fidelity: While reference baselines like SnapKV achieve high absolute scores in lenient environments, they typically exhibit a performance drop of 7–10% compared to their corresponding full-context baselines (e.g., SnapKV 30.61 vs. Ref-FullKV 32.92 on Llama-3). In contrast, within our strictly unified setting, S3-Hybrid demonstrates exceptional fidelity. For Llama-3-8B, it achieves an average score of 24.87, retaining 99.4% of the internal FullKV performance (25.01).
  • Stability Across Models: This trend holds for Qwen2-7B, where S3-Hybrid (18.80) matches the FullKV baseline (18.92) within the margin of error (>>99% retention), significantly outperforming the sparse retrieval baseline BM25 (18.80 vs 18.80, note: tied in avg but S3 is higher in Multi-Doc). …demonstrating that our SAE-driven indexing introduces negligible information loss compared to the theoretical upper bound.
The ”Denoising” Effect in Information-Dense Tasks.
4.4 Comparison with Exogenous Retrieval (RAG)
Dataset Layers_1 Layers_2 Layers_3 Layers_4
Llama-3.1-8B-Instruct
NarrativeQA 15.56 14.67 14.88 16.53
Qasper 21.41 21.56 22.34 22.75
MultiFieldQA 45.74 45.81 45.86 45.41
HotpotQA 27.84 26.31 26.42 26.35
Mistral-7B-Instruct-v0.3
HotpotQA 18.73 19.00 19.16 19.26
GovReport 16.59 16.90 16.99 16.73
Qwen2-7B-Instruct
2WikiMQA 15.14 16.56 16.83 16.04
Qasper 20.35 19.01 19.44 20.21
Method NLL $\downarrow$ Recall $\uparrow$ KL $\downarrow$
S3-Hybrid (Ours) 1.8573 0.8400 0.2154
S3-Pure (Ours) 2.0652 0.7800 0.6510
BM25 1.9593 0.7700 0.3707
RAG 1.8630 0.7700 0.3831
@@ 5 Conclusion
Overall, our results suggest that attention-aligned semantic indexing is a promising direction for memory-bounded long-context inference, with substantial room for systems optimization to reach competitive latency.
@@ B.3 Hyperparameters & Training Configuration
Baseline Configuration. For the RAG baseline, we use BAAI/bge-small-en-v1.5 as the embedding model. The text is chunked into 256-token segments with no overlap. We retrieve the top-$k$ chunks such that the total retrieved token count matches the budget of S3-Attention (approx. 1024-2048 tokens) to ensure a fair comparison.
@@ Appendix D Theoretical Analysis
D.1 From Information Bottleneck to Feature Matching
We provide a formal justification for why SAE feature co-occurrence approximates mutual information maximization.
Assumption D.1 (Sparse Disentanglement).
The SAE latent space is approximately disentangled: each feature $f$ encodes an independent semantic concept, and the joint distribution factorizes as $P(\mathcal{F})\approx\prod_{f}P(f)$.
Proposition D.2 (Proposition (Motivating bound under simplifying assumptions)).
Under Assumption D.1, let $\mathcal{F}_{Q}$ denote the active features when encoding the query $Q$, and $\mathcal{F}_{C}=\bigcup_{t\in C}\mathcal{F}_{t}$ the features present in the full context. For a subset $\hat{C}\subseteq C$, we have:
@@ D.1 From Information Bottleneck to Feature Matching
$I(Y;\hat{C}\mid Q)\geq\alpha\cdot|\mathcal{F}_{Q}\cap\mathcal{F}_{\hat{C}}|-\text{const}$ (9)
where $\alpha>0$ depends on the feature informativeness and $\mathcal{F}_{\hat{C}}=\bigcup_{t\in\hat{C}}\mathcal{F}_{t}$.

### v2

@@ Abstract
Long-context inference in Large Language Models (LLMs) faces a critical dilemma: maintaining full context incurs linear KV cache scaling, while offloading to external retrievers often yields lexically similar but causally irrelevant passages. To bridge this gap, we present $S^{3}$-Attention, a framework that transforms memory-bound inference into a streaming, attention-aligned endogenous retrieval process. Our approach distinguishes itself by decoding transient attention states into Top-$k$ sparse feature IDs via lightweight sparse autoencoders. Instead of maintaining a massive GPU Key-Value cache, we build a CPU-based inverted index during a streaming scan, ensuring GPU memory remains constant and bounded by chunk size. This mechanism aligns retrieval directly with the model’s inherent reasoning patterns, using feature co-activation (optionally fused with BM25) to recall compact evidence spans. Empirically, under the unified LongBench protocol, $S^{3}$-Hybrid closely matches full-context performance across multiple model families and improves robustness in information-dense settings by effectively filtering noise.
@@ 1 Introduction
Large language models (LLMs) have evolved into the default interface for complex cognitive tasks, transitioning from processing single documents to digesting massive collections of reports, codebases, and conversational histories. This shift toward contexts that far exceed a single attention window is not merely a technical scaling challenge but a fundamental requirement for the next generation of AI systems (Ding et al., 2024; Zhao et al., 2024a). However, bridging the gap between the theoretical capability of processing million-token regimes and the practical reality of efficient deployment remains a critical bottleneck.
@@
The prevailing response has been to extend context lengths through continued training or architectural modifications (Liu et al., 2025b; Chen et al., 2024b; Hu et al., 2024; Du et al., 2025; Mao et al., 2025; Mohtashami and Jaggi, 2023). Yet, longer windows do not automatically translate into reliable reasoning. Full-context inference is deployment-unfriendly (Ma et al., 2024): self-attention incurs quadratic compute costs (Lou et al., 2024), and Key-Value (KV) caches scale linearly (Sun et al., 2024) with input length (Tang et al., 2024), quickly saturating GPU memory (Wang et al., 2025a). Furthermore, real-world long inputs are typically sparse in signal—only a small fraction of tokens are causally useful (Zhu et al., 2025)—so naively attending to everything often dilutes evidence and amplifies distraction (Xu et al., 2024; Hooper et al., 2025; Zarch et al., 2025). The pragmatic alternative, Retrieval-Augmented Generation (RAG) (Lewis et al., 2020; Cheng et al., 2025), solves the memory issue but introduces a “semantic mismatch.” Because external retrievers operate in an embedding space independent of the generator’s internal reasoning features (Wei et al., 2025), they often retrieve lexically similar but causally irrelevant text, degrading multi-hop reasoning (Liu et al., 2025a) and increasing hallucinations (Wang et al., 2025b).
This dilemma prompts a fundamental question: Can we achieve the memory efficiency of RAG while retaining the cognitive alignment of full-context attention? This motivates a shift toward endogenous retrieval, where the model retrieves evidence using its own internal signals (Wu et al., 2024; Zhang et al., 2025). While recent methods like InfiniRetri (Ye et al., 2025) and various KV-compression techniques (Xiao et al., 2024a) utilize attention patterns to locate information, they lack a scalable mechanism for indexing. Directly operating on dense attention weights or cached states does not yield a viable queryable memory (Chen et al., 2025b): these continuous, high-dimensional representations are too expensive to store or search at token granularity (Ma et al., 2024; Liu et al., 2024, 2025c). Existing systems effectively compress memory but fail to provide an explicit, searchable index that allows for precise evidence retrieval.
@@
To realize this vision, we propose S3-Attention (Sparse & Semantic Streaming Attention), a framework designed specifically for settings where GPU memory is the primary bottleneck and causal evidence is sparse. S3-Attention transforms long-context inference into a streaming, attention-aligned retrieval procedure. The core challenge is converting the model’s dense internal states into a format that is both lightweight and efficiently searchable. We address this by training lightweight Top-$k$ sparse autoencoders (SAEs) to compress the transient key and query projections—which are already computed by the model—into discrete sparse semantic features. During a single streaming prefill pass, S3-Attention builds an inverted index on the CPU and immediately discards the KV cache, effectively achieving $O(1)$ GPU memory with respect to total context length. At generation time, the query’s SAE features retrieve high-density spans via feature co-activation. To ensure robustness against rare entities that may elude semantic compression, we optionally fuse this endogenous signal with BM25 lexical matching, yielding S3-Hybrid.
@@
Empirically, S3-Attention demonstrates exceptional fidelity on the LongBench suite. S3-Hybrid retains 99.4% of full-context performance on Llama-3-8B (24.87 vs. 25.01) and over 99% on Qwen2-7B, while enabling constant-GPU-memory processing. Notably, we observe a “denoising” effect on information-dense tasks, where our selective processing filters distraction and occasionally outperforms full-context baselines.
@@
  • We articulate long-context inference as an endogenous retrieval problem and motivate why aligning retrieval with internal attention representations mitigates the semantic gap of exogenous retrievers.
  • We introduce S3-Attention, which utilizes SAE-decoded sparse semantic features to build a streaming inverted index, achieving $O(1)$ GPU memory without fine-tuning the base LLM.
  • We demonstrate that fusing endogenous semantic signals with BM25 yields a robust hybrid retriever with near-lossless fidelity to full-context inference, validated across multiple model families on LongBench.
@@ 3.2 Theoretical Framework: The Endogenous Information Bottleneck
Instead, we propose an operational endogenous proxy derived from the model’s own attention-matching signals. Specifically, we discretize transient attention projections into a sparse feature space and rank context positions by their inverse-document-frequency (IDF) weighted feature co-activation with the query. Appendix D provides a motivating inequality under explicit simplifying assumptions. We emphasize that this result should be interpreted as a heuristic justification of our scoring rule, rather than as a guarantee that feature overlap faithfully tracks mutual information in real models.
@@ 4.2 Qualitative Analysis: Mechanism of Endogenous Alignment
Figure 2: Endogenous vs. Exogenous Retrieval. Top: RAG (BGE-Small) is distracted by high lexical overlap… Bottom: S3-Attention (Ours) ignores the distraction… (For a larger view, please refer to Figure 4 in the Appendix.)
Additional visualization results and analysis of other datasets are provided in Appendix G.
@@
4.3 Quantitative Results on LongBench
We present the comprehensive evaluation results on LongBench in Table 1. It is important to note that absolute evaluation scores on LongBench can fluctuate significantly depending on inference environments (e.g., prompt templates and quantization kernels). To ensure a rigorous comparison, we distinguish between standard baselines (provided for reference) and methods evaluated within our unified environment (marked in bold), which include FullKV, RAG(with rerank), BM25, and our S3-Attention variants.We also present an additional analysis of $S^{3}$ versus retrieval-based methods in the zero-shot setting; see H.
Performance Retention: Near-Lossless Compression.
  • Superior Fidelity: While reference baselines like SnapKV achieve high absolute scores in lenient environments, they typically exhibit a performance drop of 7–10% compared to their corresponding full-context baselines (e.g., SnapKV 42.96 vs. Ref-FullKV 49.74 on Llama-3.1). In contrast, within our strictly unified setting, S3-Hybrid demonstrates exceptional fidelity. For Llama-3-8B, it achieves an average score of 24.87, retaining 99.4% of the internal FullKV performance (25.01).
  • Stability Across Models: This trend holds for Mistral-7B, where S3-Hybrid (23.24) matches the FullKV baseline (23.40) within the margin of error (>>99% retention), better than the sparse retrieval baseline BM25 (23.24 vs 22.72) ,demonstrating that our SAE-driven indexing introduces negligible information loss compared to the theoretical upper bound.
The ”Denoising” Effect in Information-Dense Tasks.
4.4 Comparison with Exogenous Retrieval (RAG)
Dataset Layers_1 Layers_2 Layers_3 Layers_4
Llama-3.1-8B-Instruct
NarrativeQA 15.56 14.67 14.88 16.53
Qasper 21.41 21.56 22.34 22.75
MultiFieldQA 45.74 45.81 45.86 45.41
HotpotQA 27.84 26.31 26.42 26.35
Mistral-7B-Instruct-v0.3
HotpotQA 18.73 19.00 19.16 19.26
GovReport 16.59 16.90 16.99 16.73
Qwen2-7B-Instruct
2WikiMQA 15.14 16.56 16.83 16.04
Qasper 20.35 19.01 19.44 20.21
Method NLL $\downarrow$ Recall $\uparrow$ KL $\downarrow$
S3-Hybrid (Ours) 1.8573 0.8400 0.2154
S3-Pure (Ours) 2.0652 0.7800 0.6510
BM25 1.9593 0.7700 0.3707
RAG 1.8630 0.7700 0.3831
@@ 5 Conclusion
Overall, our results suggest that attention-aligned semantic indexing is a promising direction for memory-bounded long-context inference, with substantial room for systems optimization to reach competitive latency. We also report the current engineering limitation that our prototype incurs higher wall-clock latency than optimized FullKV baselines, motivating future kernel-level optimization.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ B.3 Hyperparameters & Training Configuration
Baseline Configuration. To ensure a fair and rigorous comparison, we implemented a unified benchmarking framework that explicitly measures wall-clock latency. For the RAG baseline, we adopt a robust two-stage retrieval strategy utilizing BAAI/bge-small-en-v1.5 as the embedding model. The corpus is processed into non-overlapping segments of 256 tokens. During retrieval, the top-100 candidates are first identified via the dense retriever; these candidates are subsequently re-scored by a cross-encoder (BGE-Reranker-v2-m3) to select the final top-40 chunks. This configuration is designed to align the total retrieved token count with the computational budget of S3-Attention (approximately 1024–2048 tokens).
@@ Appendix D Theoretical Analysis
D.1 Preliminaries and Notation
We first establish precise notation aligned with our implementation:
  • $C=\{c_{1},c_{2},\ldots,c_{L}\}$: Context token sequence of length $L$
  • $Q$: Query representation
  • $\mathcal{F}_{t}\subseteq[D]$: Set of $k$ active SAE features at position $t$, where $|\mathcal{F}_{t}|=k$
  • $a_{f}^{(t)}\geq 0$: Activation magnitude of feature $f$ at position $t$
  • $\text{freq}(f)=|\{t:f\in\mathcal{F}_{t}\}|$: Document frequency of feature $f$
D.2 The Scoring Function: A Pragmatic Derivation
Rather than claiming a formal information-theoretic bound, we provide a principled heuristic justification for our scoring function based on retrieval theory.
D.2.1 The Implemented Scoring Function
Our implementation computes the relevance score for each context position $t$ as:
$s(t)=\sum_{f\in\mathcal{F}_{Q}}a_{f}^{(Q)}\cdot\text{IDF}(f)\cdot\mathbb{1}[f\in\mathcal{F}_{t}]$ (9)
where:
$\text{IDF}(f)=\frac{1}{\log(1+\text{freq}(f))+1}$ (10)
D.2.2 Justification via Weighted Feature Matching
Proposition D.1 (Scoring as Weighted Jaccard Similarity).
The scoring function in Eq. 9 can be interpreted as a weighted soft Jaccard similarity:
$s(t)\propto\sum_{f\in\mathcal{F}_{Q}\cap\mathcal{F}_{t}}w_{f}^{(Q)}$ (11)
where $w_{f}^{(Q)}=a_{f}^{(Q)}\cdot\text{IDF}(f)$ assigns higher weight to:
  1. 1.
    Features with strong query activation ($a_{f}^{(Q)}$ large)
  2. 2.
    Features that are rare in the context ($\text{freq}(f)$ small)
D.3 Information-Theoretic Motivation (Informal)
We provide an informal motivation connecting feature matching to information preservation, without claiming a rigorous bound.
D.3.1 The Compression-Utility Tradeoff
In the context compression setting, we seek a subset $\hat{C}\subseteq C$ that:
  1. 1.
    Preserves utility: Retains information relevant to answering $Q$
  2. 2.
    Achieves compression: $|\hat{C}|\ll|C|$
Remark D.2 (Connection to Mutual Information).
Under the informal assumption that SAE features capture semantically meaningful concepts, positions $t$ where $\mathcal{F}_{t}\cap\mathcal{F}_{Q}\neq\emptyset$ are more likely to contain information relevant to the query. We do not claim this maximizes $I(Y;\hat{C}\mid Q)$ in a formal sense.
D.3.2 Why Feature Overlap is a Reasonable Proxy
SAE features trained with sparsity constraints tend to capture interpretable semantic concepts. When query features overlap with context features:
  • The context position likely discusses concepts mentioned in the query
  • Such positions have higher probability of containing answer-relevant information
This is an empirical observation, not a mathematical theorem.
D.4 IDF Weighting: Precision Enhancement
D.4.1 The Role of IDF in Retrieval
The IDF weighting serves a well-understood purpose in information retrieval:
Proposition D.3 (IDF as Discriminative Weighting).
Features with high document frequency provide less discriminative power for retrieval:
$\text{Discriminative Power}(f)\propto\frac{1}{\text{freq}(f)}$ (12)
The IDF formula $\text{IDF}(f)=\frac{1}{\log(1+\text{freq}(f))+1}$ implements a smoothed version of this principle.
D.4.2 Clarification on the Original IDF Justification
The original statement (“Reduces contribution of high-frequency features to $I(Y;\hat{C}\mid Q)$ without reducing I(C;C^)I($\hat{C}$)”) was imprecise. A clearer formulation:
IDF weighting increases retrieval precision by down-weighting features that match many positions indiscriminately. This focuses the retrieval budget on positions that share distinctive features with the query, rather than common features that provide little signal about relevance.
D.5 Corrected Theoretical Statement
We replace the original Proposition with a more honest characterization:
Proposition D.4 (Feature Matching Heuristic).
Let $s(t)$ be computed according to Eq. 9, and let $\hat{C}=\{t:s(t)\geq\tau\}$ be the retrieved subset. Under the following empirical assumptions:
  1. (A1)
    SAE features correspond to interpretable semantic concepts
  2. (A2)
    Query-relevant context positions activate similar features to the query
  3. (A3)
    High-frequency features are less informative for relevance discrimination
The scoring function provides a computationally efficient proxy for identifying query-relevant context positions.
Remark D.5 (On the Original Proof).
@@ D.5 Corrected Theoretical Statement
The original proof attempted to establish $I(Y;\hat{C}\mid Q)\geq\alpha\cdot|\mathcal{F}_{Q}\cap\mathcal{F}_{\hat{C}}|-\text{const}$ through:
  • Step 2 (DPI): The claimed inequality $I(Y;\hat{C}\mid Q)\geq I(Y;\mathcal{F}_{\hat{C}}\mid\mathcal{F}_{Q})$ requires a Markov chain $Y-\hat{C}-\mathcal{F}_{\hat{C}}$ given $Q$, which was not established.
  • Step 3 (Decomposition): The additive decomposition of mutual information requires stronger independence assumptions than marginal independence of features.
  • Step 4 (Bound): The definition $\alpha=\min_{f\in\mathcal{F}_{Q}}I(Y;f\mid\mathcal{F}_{Q})$ may equal zero if feature presence is determined by $\mathcal{F}_{Q}$.
We acknowledge these issues and instead provide the pragmatic justification above.
D.6 Complexity Analysis
Proposition D.6 (Memory Complexity).
The indexing phase achieves $\mathcal{O}(1)$ GPU memory complexity with respect to context length $L$.